### v2

# Uniform Set Systems with Uniform Witnesses

Abstract.
Frankl and Pach [undefg] and Erdős [undefd] conjectured that any $(d+1)$-uniform set family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ with VC-dimension at most $d$ has size at most $\binom{n-1}{d}$ when $n$ is sufficiently large. Ahlswede and Khachatrian [undef] showed that the conjecture is false by giving a counterexample of size $\binom{n-1}{d}+\binom{n-4}{d-2}$.
For a set family $\mathcal{F}\subseteq\binom{[n]}{d+1}$, the condition that its VC-dimension is at most $d$ can be reformulated as follows: for any $F\in\mathcal{F}$, there exists a set $B_{F}\subseteq F$ such that $F\cap F^{\prime}\neq B_{F}$ for all $F^{\prime}\in\mathcal{F}$. In this direction, the first author, Xu, Yip, and Zhang [undefc] conjectured that the bound $\binom{n-1}{d}$ holds if we further assume that $|B_{F}|=s$ for every $F\in\mathcal{F}$ and for some fixed $0\leqslant s\leqslant d$.
The case $s=0$ is exactly the Erdős–Ko–Rado theorem, and the cases $s\in\{1,d\}$ were proved in the paper by the first author, Xu, Yip, and Zhang [undefc]. In this short note, we show that the conjecture holds when $s\leqslant d/2$, and the maximal constructions are stars. Moreover, we construct non-star set families of size $\binom{n-1}{d}$ satisfying the condition for $d/2<s\leqslant d-1$, which suggests that the problem is substantially different in these cases.
*Department of Mathematics, Massachusetts Institute of Technology, Cambridge, MA, USA. Email: {twchao, zixuanxu,zakhdm}@mit.edu
1. Introduction
Let $\mathcal{F}\subseteq\binom{[n]}{d+1}$ be a $(d+1)$-uniform set system over the ground set $n]\overset{\rm{def}}{=}\{1,\dots,n\}$. The Vapnik-Chervonenkis dimension (VC-dimension) of $\mathcal{F}$, denoted as $\textup{VC}(\mathcal{F})$ is defined to be the size of the largest set $S\subseteq[n]$ such that every subset of $S$ is realized as an intersection $S\cap F$ for some $F\in\mathcal{F}$, i.e. for all $S^{\prime}\subseteq S$, there exists $F\in\mathcal{F}$ such that $S^{\prime}=S\cap F$. Since every set family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ has VC-dimension at most $d+1$, it is natural to ask for the maximum size of $\mathcal{F}$ with VC-dimension at most $d$. Frankl and Pach [undefg] proved that any such set family has size at most $\binom{n}{d}$. Then Frankl and Pach [undefg] and Erdős [undefd] conjectured that the upper bound can be improved to $\binom{n-1}{d}$ when $n$ is large enough. However, Ahlswede and Khachatrian constructed a counterexample of size $\binom{n-1}{d}+\binom{n-4}{d-2}$, and this was later generalized by Mubayi and Zhao [undefi] where they constructed a class of constructions attaining the same size. The upper bound was recently improved to $\binom{n-1}{d}+O_{d}(n^{d-2})$ by Yang and Yu [undefk], matching the second order term of the family constructed in [undefi] up to a multiplicative constant. For more works along this line, see [undefc, undefh, undefj].
Notice that in the case when $\mathcal{F}$ is a $(d+1)$-uniform family, we can equivalently formulate the condition that $\textup{VC}(\mathcal{F})\leqslant d$ as follows: for each $F\in\mathcal{F}$, there exists a subset $B_{F}\subseteq F$ such that for any $F^{\prime}\in\mathcal{F}$, we have $F\cap F^{\prime}\neq B_{F}$. We call $B_{F}$ a witness of $F$ in the sense that it cannot be realized as an intersection of $F$ with any set $F^{\prime}\in\mathcal{F}$. In this note, we are interested in the case where all witness sets $B_{F}$ have the same size.
Definition 1.1 ($s$-witness family).
Let $n\geqslant d+1$ and $0\leqslant s\leqslant d$. A family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ is an $s$-witness family if it satisfies the following condition: for every $F\in\mathcal{F}$, there exists $B_{F}\subseteq F$ of size $s$ such that $F\cap F^{\prime}\neq B_{F}$ for every $F^{\prime}\in\mathcal{F}$.
A 0-witness family $\mathcal{F}$ is exactly an intersecting family since we have $F\cap F^{\prime}\neq\varnothing$ for all $F,F^{\prime}\in\mathcal{F}$. The well-known Erdős–Ko–Rado Theorem states the following upper bound for intersecting families.
Theorem 1.2 (Erdős–Ko–Rado [undefe]).
Let $d\geqslant 0$ and $n\geqslant 2(d+1)$. If $\mathcal{F}\subseteq\binom{[n]}{d+1}$ is an intersecting family, then $|\mathcal{F}|\leqslant\binom{n-1}{d}$. Moreover, when $n>2(d+1)$, equality holds if and only if $\mathcal{F}$ is a star.
As a generalization of the Erdős–Ko–Rado Theorem, the first author, Xu, Yip, and Zhang made the following conjecture in [undefc].
Conjecture 1.3 ([undefc]).
Let $n\geqslant 2(d+1)$ and $0\leqslant s\leqslant d$. Suppose $\mathcal{F}\subseteq\binom{[n]}{d+1}$ is an $s$-witness family. Then $|\mathcal{F}|\leqslant\binom{n-1}{d}$.
It is clear that when $s=0$, 1.3 follows from Theorem 1.2. In [undefc], the authors confirmed 1.3 in the case where $s\in\{1,d\}$ and $n$ is sufficiently large. Moreover, they showed that equality holds if and only if $\mathcal{F}$ is a star when $s\in\{1,d\}$ and remarked that it might be plausible to conjecture that the maximum set families achieving the bound in 1.3 must be stars for all $0\leqslant s\leqslant d$.
In this note, we prove 1.3 in the case when $s\leqslant d/2$ and $n$ is sufficiently large.
Theorem 1.4.
Let $1\leqslant s\leqslant d/2$ and let $n$ be sufficiently large. Suppose $\mathcal{F}\subseteq\binom{[n]}{d+1}$ is an $s$-witness family. Then $|\mathcal{F}|\leqslant\binom{n-1}{d}$. Moreover, if $|\mathcal{F}|=\binom{n-1}{d}$, then $\mathcal{F}$ is a star.
Together with Theorem 1.2 and the results in [undefc], the remaining open cases of 1.3 are when $d/2<s\leqslant d-1$. In these remaining open cases, we give a construction that is not a single star achieving the size $\binom{n-1}{d}$, illustrating the barriers to adapting current methods as all the proofs of the known cases show that $\mathcal{F}$ must be a single star if equality is achieved.
Theorem 1.5.
For any $n\geqslant s+3$ and $d/2<s\leqslant d-1$, there exists an $s$-witness family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ of size $\binom{n-1}{d}$ that is not a star.
We will prove Theorem 1.5 by giving an explicit construction that is contained in the union of two star families but not contained in a single star family. Recall that the piercing number $\tau(\mathcal{F})$ is the cardinality of the smallest set $T$ such that $F\cap T\neq\varnothing$ for all $F\in\mathcal{F}$. In particular, the construction given in the proof of Theorem 1.5 has piercing number $2$. Furthermore, we can generalize the construction in Theorem 1.5 and get $s$-witness families of size $\binom{n-1}{d}$ with arbitrary large piercing number.
Theorem 1.6.
Let $d,s,k\geqslant 1$ be integers such that $d/2<s\leqslant d-k+1$ and $n\geqslant\max(k+(k-1)s,d+2k-1)$. There exists an $s$-witness family $\mathcal{F}\subseteq{[n]\choose d+1}$ of size $n-1\choose d$} such that $\tau(\mathcal{F})=k$.
We remark that our construction gives several pairwise non-isomorphic examples (up to a permutation of the ground set) proving Theorem 1.5, but the construction for Theorem 1.6 gives only one such example (up to a permutation of the ground set). Therefore, we will present both the proof of Theorem 1.5 and the proof of Theorem 1.6 in Section 3 for completeness despite the fact that they share many similarities. [^1] In an earlier arXiv version of this paper, Theorem 1.6 was stated as a question in the special case $k=3$. ChatGPT 5.4 Pro managed to successfully answer this question and the construction in Section 3.2 is a generalization of its response (see https://chatgpt.com/share/69b8e49c-6df0-8009-8180-52161f034bc2 for the conversation).
Since our proof of Theorem 1.4 proceeds by approximating $\mathcal{F}$ by a single star, it does not generalize easily to the cases where $s>d/2$. Thus, we believe new ideas are required to prove the remaining cases of 1.3.
1.1. Paper Organization
In Section 2 we present the proof of Theorem 1.4. Then in Section 3 we give the construction that proves Theorems 1.5 and 1.6.
Notations. For sets $A,B\subseteq[n]$, we denote their symmetric difference as $A\triangle B=(A\setminus B)\cup(B\setminus A)$. We always consider $s$ and $d$ as fixed and $n\to\infty$. We use $O_{d}(\cdot)$ and $\Omega_{d}(\cdot)$ to suppress leading constants depending on $d$.
Acknowledgments. We thank the organizers of the SLMath workshop “Algebraic and Analytic Methods in Combinatorics” (March 2025) attended by the authors where this project was initiated. We would also like to thank Zixiang Xu for comments on an early draft of this note. Finally, we thank Zach Hunter for pointing out typos in the first draft posted on arXiv.
2. Proof of Theorem 1.4
Our proof consists of three steps split into the three subsequent subsections. In the first step presented in Section 2.1, we analyze the structure of an $s$-witness set family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ for $1\leqslant s\leqslant d/2$. In particular, we show in Lemma 2.2 that we can capture the intersection properties of $\mathcal{F}$ using set families of constant size. Then in Lemma 2.4, we define a certain injection of a subset of $\mathcal{F}$ into $\binom{[n]}{d}$ that will be the main object that we analyze later on. Next, in Section 2.2, we show that any family $\mathcal{F}$ that is close to achieving the maximum size can be approximated by a star family. Finally, in Section 2.3, we perform a stability analysis, showing that in fact any set family achieving size $\binom{n-1}{d}$ must be a single star, concluding the proof of Theorem 1.4.
From now on, we fix $1\leqslant s\leqslant d/2$ and let $\mathcal{F}\subseteq\binom{[n]}{d+1}$ be an $s$-witness family. In particular, for $F\in\mathcal{F}$, we always fix a choice of its witness $B_{F}\subseteq F$ of size $s$ so that $F\cap F^{\prime}\neq B_{F}$ for any $F^{\prime}\in\mathcal{F}$.
Let us recall the Erdős-Rado sunflower lemma that will be used in the proof of Lemma 2.2. For $r\geqslant 2$, a family of $r$ sets $H_{1},\dots,H_{r}\subseteq[n]$ is called a $k$-sunflower if there exists $S\subseteq[n]$ such that $H_{i}\cap H_{j}=S$ for all $i\neq j\in[r]$.
Lemma 2.1 (Erdős-Rado sunflower lemma [undeff]).
Let $k$ and $r\geqslant 2$ be positive integers. If $\mathcal{F}\subseteq\binom{[n]}{k}$ has size $|\mathcal{F}|\geqslant k!(r-1)^{k}$, then $\mathcal{F}$ contains an $r$-sunflower.
The bound in Lemma 2.1 was later improved by Alweiss-Lovette-Wu-Zhang [undefa] and Bell–Chueluecha–Warnke [undefb], but the bound stated in Lemma 2.1 suffices for our purposes. Now we begin the proof of Theorem 1.4.
2.1. Structural Analysis
In this section, we prove that one can capture the intersection properties in $\mathcal{F}$ using set families of constant size. First, observe that we can interpret the uniform witness condition as follows. For any $B\in\binom{[n]}{s}$ and $F,F^{\prime}\in\mathcal{F}$ such that $B_{F}=B$ and $B\subseteq F^{\prime}$, we must have $(F\setminus B)\cap(F^{\prime}\setminus B)\neq\varnothing$. In other words, the two families $\{F\setminus B\mid F\in\mathcal{F},\,B_{F}=B\}$ and $\{F^{\prime}\setminus B\mid F^{\prime}\in\mathcal{F},\,B\subseteq F^{\prime}\}$ are cross-intersecting.
Define $\mathcal{B}=\{B\in\binom{[n]}{s}\mid\exists F\in\mathcal{F}\text{ such that }B_{F}=B\}$ to be the set of all witnesses for $\mathcal{F}$. We have the following modeling lemma that captures for each $B\in\mathcal{B}$ the structure of the sets $F\in\mathcal{F}$ containing $B$.
Lemma 2.2 (Modeling Lemma).
For every $B\in\mathcal{B}$, there exists a set family $\mathcal{A}_{B}\subseteq\binom{[n]\setminus B}{\leqslant d+1-s}$ consisting of sets of size at most d+1sd+1-s satisfying the following:
  1. (1)
    $\mathcal{A}_{B}$ is an intersecting family, i.e. $A\cap A^{\prime}\neq\varnothing$ for all $A,A^{\prime}\in\mathcal{A}_{B}$.
  2. (2)
    For every $F\in\mathcal{F}$ such that $B_{F}=B$, there exists $A\in\mathcal{A}_{B}$ such that $A\subseteq F$.
  3. (3)
    For every $F\in\mathcal{F}$ such that $B\subseteq F$ (possibly $B_{F}\neq B$), we have $A\cap F\neq\varnothing$ for all $A\in\mathcal{A}_{B}$.
  4. (4)
    $|\mathcal{A}_{B}|=O_{d}(1)$.
Proof.
We construct $\mathcal{A}_{B}$ iteratively. Initialize $\mathcal{A}\overset{\rm{def}}{=}\{F\setminus B\mid F\in\mathcal{F},B_{F}=B\}$. If $\mathcal{A}$ contains a $(d+2)$-sunflower $A_{1},\dots,A_{d+2}\in\mathcal{A}$ with a core $S\subseteq[n]\setminus B$, then we remove the sets $A_{1},\dots,A_{d+2}$ from $\mathcal{A}$ and add the set $S$ to $\mathcal{A}$. We repeat this operation until $\mathcal{A}$ is $(d+2)$-sunflower-free and set $\mathcal{A}_{B}$ as the final ($d+2$)-sunflower-free family. We show that $\mathcal{A}_{B}$ satisfies all the properties in Lemma 2.2.
First, note that the initial family $\mathcal{A}_{0}=\{F\setminus B\mid F\in\mathcal{F},B_{F}=B\}$ satisfies properties (1)-(3). Indeed, $\mathcal{A}_{0}$ must be an intersecting family, because if there exists $A\cap A^{\prime}=\varnothing$, then the sets $(A\cup B),(A^{\prime}\cup B)\in\mathcal{F}$ violate the property that $B$ is the witness of $A\cup B$ and $A^{\prime}\cup B$. Property (2) is trivially satisfied by definition. Property (3) also follows from the fact that $B$ is the witness for sets of the form $A\cup B$ where $A\in\mathcal{A}_{0}$. For a set $F\supseteq B$, we must have $F\cap(A\cup B)\neq B$ and thus $F\cap A\neq\varnothing$ for all $A\in\mathcal{A}_{0}$.
[…]
Finally, we upper bound the size of $\mathcal{A}_{B}$. The family $\mathcal{A}_{B}$ is ($d+2$)-sunflower-free and consists of sets of size at most d+1sd+1-s. We must have $\mathcal{A}_{B}$ is ($d+2$)-sunflower-free at every uniformity. Thus by Lemma 2.1 we have
$|\mathcal{A}_{B}|\leqslant\sum_{k=1}^{d+1-s}k!(d+1)^{k}=O_{d}(1).$
This completes the proof. ∎
Define
[…]
In particular, we have $F\in\mathcal{F}_{B}$ if $B_{F}=B$, but $\mathcal{F}_{B}$ may also include other sets. Note that at this point we no longer need to keep track of the particular choice of $B_{F}$ for each $F\in\mathcal{F}$, instead we record that
$\mathcal{F}=\bigcup_{B\in\mathcal{B}}\mathcal{F}_{B}.$
For $B\in\mathcal{B}$, define $\alpha_{B}=\min_{A\in\mathcal{A}_{B}}|A|$.
Claim 2.3.
For any $B\in\mathcal{B}$, we have
$|\mathcal{F}_{B}|=O_{d}\left(n^{d+1-s-\alpha_{B}}\right).$
Proof.
Since each $F\in\mathcal{F}_{B}$ contains an element of $\mathcal{A}_{B}$ as a subset, we get by double counting and the definition of $\alpha_{B}$:
$|\mathcal{F}_{B}|\leqslant|\mathcal{A}_{B}|\cdot{n-s-\alpha_{B}\choose d+1-s-|A|}=|\mathcal{A}_{B}|\cdot O_{d}\left(n^{d+1-s-\alpha_{B}}\right).$
By Item 4 in Lemma 2.2, we have $|\mathcal{A}_{B}|=O_{d}(1)$ and thus we have the desired bound. ∎
Define $\mathcal{B}_{1}=\{B\in\mathcal{B}\mid\alpha_{B}=1\}$ and $\mathcal{B}_{\geqslant 2}=\mathcal{B}\setminus\mathcal{B}_{1}$. For $B\in\mathcal{B}_{1}$ let $x_{B}$ be the unique element so that $\{x_{B}\}\in\mathcal{A}_{B}$ ($x_{B}$ is unique since $\mathcal{A}_{B}$ is an intersecting family). Since $\mathcal{A}_{B}$ is intersecting, if $\{x_{B}\}\in\mathcal{A}_{B}$, then we can redefine $\mathcal{A}_{B}=\{\{x_{B}\}\}$ and the properties in Lemma 2.2 are still satisfied. Now we define an injection of a subset of $\mathcal{F}$ into $\binom{[n]}{d}$ which will be a crucial part of our proof.
Lemma 2.4 (Injective mapping $F\rightarrow E$).
For $B\in\mathcal{B}_{1}$, define $\mathcal{E}_{B}=\{F\setminus\{x_{B}\}\mid F\in\mathcal{F}_{B}\}$ and let $\mathcal{E}=\bigcup_{B\in\mathcal{B}_{1}}\mathcal{E}_{B}$. Then we have the following properties:
  1. (1)
    If $F\in\mathcal{F}$, $B\in\mathcal{B}_{1}$ and $E\in\mathcal{E}_{B}$ are such that $E\subseteq F$, then we have $F=E\cup\{x_{B}\}$.
  2. (2)
    For $B\neq B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$, we have $\mathcal{E}_{B}\cap\mathcal{E}_{B^{\prime}}=\varnothing$.
  3. (3)
    For $B,B^{\prime}\in\mathcal{B}_{1}$, let
    $\mathcal{U}(B,B^{\prime})=\left\{E\in{[n]\choose d}\,\middle|\,B\cup B^{\prime}\subseteq E,~x_{B},x_{B^{\prime}}\not\in E\right\}.$
    Then we have $\mathcal{U}(B,B^{\prime})\cap\mathcal{E}=\varnothing$ for any $B,B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$.
The idea of constructing an injection from $\mathcal{F}$ into $n]\choose d$} appears in earlier works as well (see [undefc, Claim 3.1] and [undefk]). However, our implementation of this idea is different and, in particular, Item 3 in Lemma 2.4 is new.
Proof.
We first prove Item 1. From the assumption, we have $B\subseteq E\subseteq F$. By Item 3 in Lemma 2.2, we know that $F\cap A\neq\varnothing$ for all $A\in\mathcal{A}_{B}$. Since $B\in\mathcal{B}_{1}$, we have $\mathcal{A}_{B}=\{\{x_{B}\}\}$. Thus, $x_{B}\in F$ and we must have $F=E\cup\{x_{B}\}$.
To prove Item 2, suppose for contradiction that we have $E\in\mathcal{E}_{B}\cap\mathcal{E}_{B^{\prime}}$ for $B$ and $B^{\prime}$ with $x_{B}\neq x_{B^{\prime}}$. Then we have $E=F\setminus\{x_{B}\}=F^{\prime}\setminus\{x_{B^{\prime}}\}$ for some $F,F^{\prime}\in\mathcal{F}$. By Item 1 we must have $F=E\cup\{x_{B}\}=E\cup\{x_{B^{\prime}}\}$ and thus $x_{B}=x_{B^{\prime}}$, contradicting the assumption.
Finally, to prove Item 3, consider $B,B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$. Assume for contradiction that there exists $E\in\mathcal{U}(B,B^{\prime})\cap\mathcal{E}$, then we must have $E=F^{\prime\prime}\setminus\{x_{B^{\prime\prime}}\}$ for some $B^{\prime\prime}\subseteq F^{\prime\prime}\in\mathcal{F}$. On the other hand, since $B\subseteq F^{\prime\prime}$, by Item 1 we must have $x_{B^{\prime\prime}}=x_{B}$. Similarly, since $B^{\prime}\subseteq F^{\prime\prime}$, we have $x_{B^{\prime}}=x_{B}$. Therefore we have $x_{B}=x_{B^{\prime\prime}}$ contradicting the assumption that $x_{B}\neq x_{B^{\prime}}$. ∎
Define $\mathcal{F}_{1}=\bigcup_{B\in\mathcal{B}_{1}}\mathcal{F}_{B}$ and $\mathcal{F}_{\geqslant 2}=\bigcup_{B\in\mathcal{B}_{2}}\mathcal{F}_{B}$. Note that by definition $|\mathcal{F}|\leqslant|\mathcal{F}_{1}|+|\mathcal{F}_{\geqslant 2}|$. Then Lemma 2.4 implies that the map $\varphi:\mathcal{F}_{1}\to\binom{[n]}{d}$ defined by $F\mapsto F\setminus\{x_{B}\}$ is an injection. In particular, $\mathcal{E}$ is the image $\varphi(\mathcal{F})\subseteq\binom{[n]}{d}$, so we have $|\mathcal{F}_{1}|=|\mathcal{E}|$, and it suffices to upper bound $\mathcal{E}$. From Item 3 in Lemma 2.4, the set of $\mathcal{U}(B,B^{\prime})$ for $B,B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$ is contained in the complement of $\mathcal{E}$.
In order to have $\mathcal{U}(B,B^{\prime})\neq\varnothing$, we need $x_{B}\notin B^{\prime}$ and $x_{B^{\prime}}\notin B$. Thus, we define
$\mathcal{P}\overset{\rm{def}}{=}\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B}\neq x_{B^{\prime}},x_{B}\notin B^{\prime},x_{B^{\prime}}\notin B\}\quad\text{and}\quad\mathcal{U}\overset{\rm{def}}{=}\bigcup_{(B,B^{\prime})\in\mathcal{P}}\mathcal{U}(B,B^{\prime}).$
Then we have the following useful inequality for bounding $|\mathcal{F}_{1}|$ as a direct corollary of Lemma 2.4.
Corollary 2.5.
We have
$|\mathcal{F}_{1}|+|\mathcal{U}|\leqslant{n\choose d}$. (1)
2.2. Star approximation
In this section, we prove that any $s$-witness family with $1\leqslant s\leqslant d/2$ and size close to maximal can be approximated by a star family up to a bounded symmetric difference.
Recall that $\mathcal{F}_{1}=\bigcup_{B\in\mathcal{B}_{1}}\mathcal{F}_{B}$, $\mathcal{F}_{\geqslant 2}=\bigcup_{B\in\mathcal{B}_{\geqslant 2}}\mathcal{F}_{B}$, and $|\mathcal{F}|\leqslant|\mathcal{F}_{1}|+|\mathcal{F}_{\geqslant 2}|$. We have the following estimate.
Claim 2.6.
We have $|\mathcal{F}_{\geqslant 2}|=O_{d}(n^{d-1})$.
Proof.
Note that we have
$|\mathcal{F}_{\geqslant 2}|\leqslant\sum_{B\in\mathcal{B}_{\geqslant 2}}|\mathcal{F}_{B}|.$
For each $B\in\mathcal{B}_{\geqslant 2}$, we have $\alpha_{B}\geqslant 2$ and thus by Claim 2.3 we have $|\mathcal{F}_{B}|=O_{d}(n^{d-s-1})$. Therefore, we obtain
$|\mathcal{F}_{\geqslant 2}|\leqslant|\mathcal{B}_{\geqslant 2}|\cdot O_{d}(n^{d-s-1})\leqslant\binom{n}{s}\cdot O_{d}(n^{d-s-1})=O_{d}(n^{d-1}).\qed$
Now we state our result on star approximation.
Proposition 2.7.
Suppose $d\geqslant 2s$ and $n\geqslant 2(d+1)$. For any $K\geqslant 0$, if $\mathcal{F}\subseteq\binom{[n]}{d+1}$ is an $s$-witness family with size
$|\mathcal{F}|\geqslant{n-1\choose d}-Kn^{d-1},$
then there exists a star family $\mathcal{F}^{*}$ with $|\mathcal{F}\triangle\mathcal{F}^{*}|\leqslant O_{d}((K+1)n^{d-1})$.
Proof.
Note that by adjusting the implied constant in $O_{d}$, we may assume that $K\leqslant cn$ for an arbitrarily small fixed constant $c>0$ depending on $d$.
First note that by Claim 2.6, we have $|\mathcal{F}_{\geqslant 2}|=O_{d}(n^{d-1})$, so we may focus on $\mathcal{F}_{1}$. We first prove a lower bound on $|\mathcal{B}_{1}|$ in the following claim.
Claim 2.8.
We have $|\mathcal{B}_{1}|=\Omega_{d}(n^{s})$.
Proof.
From Claim 2.3, we have
$|\mathcal{F}_{1}|\leqslant\sum_{B\in\mathcal{B}_{1}}|\mathcal{F}_{B}|\leqslant|\mathcal{B}_{1}|\cdot O_{d}(n^{d-s}).$
By rearranging, we get $|\mathcal{B}_{1}|=\Omega_{d}\left(|\mathcal{F}_{1}|n^{s-d}\right)$. Recall that we have $|\mathcal{F}_{1}|\geqslant|\mathcal{F}|-|\mathcal{F}_{\geqslant 2}|$, $|\mathcal{F}_{2}|=O_{d}(n^{d-1})$, and $|\mathcal{F}|\geqslant\binom{n-1}{d}-Kn^{d-1}=\Omega_{d}(n^{d})$, where we use the assumption $K\leqslant cn$. It follows that $|\mathcal{F}_{1}|=\Omega_{d}(n^{d})$, and hence $|\mathcal{B}_{1}|=\Omega_{d}(n^{s})$. ∎
For $x\in[n]$, let $\mathcal{B}_{1}(x)\overset{\rm{def}}{=}\{B\in\mathcal{B}_{1}\mid x_{B}=x\}$. Define $x_{0}\overset{\rm{def}}{=}\arg\max_{x\in[n]}|\mathcal{B}_{1}(x)|$ and let $M\geqslant 0$ be such that $|\mathcal{B}_{1}(x_{0})|=|\mathcal{B}_{1}|-M$. Define the star family
$\mathcal{F}^{*}=\left\{F^{*}\in\binom{[n]}{d+1}\,\middle|\,x_{0}\in F^{*}\right\}.$
Note that we have
[…]
By assumption, we have
[…]
Thus it suffices to bound $|\mathcal{F}\setminus\mathcal{F}^{*}|$. By definition and Claim 2.3, we have
$|\mathcal{F}\setminus\mathcal{F}^{*}|\leqslant|\mathcal{F}_{\geqslant 2}|+\sum_{B\in\mathcal{B}_{1}\setminus\mathcal{B}_{1}(x_{0})}|\mathcal{F}_{B}|\leqslant O_{d}(n^{d-1})+M\cdot O(n^{d-s}).$
So now our goal is to upper bound $M$. We first bound it in terms of $\left\lvert\mathcal{U}\right\rvert$.
Claim 2.9.
We have
$\left\lvert\mathcal{U}\right\rvert=\Omega_{d}\left(n^{d-2s}\left\lvert\mathcal{B}_{1}\right\rvert\left(M-2\binom{n-1}{s-1}\right)\right)$.
Proof.
For any $(B,B^{\prime})\in\mathcal{P}$, since $s\leqslant d/2$, we have $|\mathcal{U}(B,B^{\prime})|\geqslant\binom{n-2s-2}{d-2s}$. Note that every $E\in{[n]\choose d}$ can contain $B\cup B^{\prime}$ for at most $\binom{d}{s}\cdot\binom{d-s}{s}$ pairs $(B,B^{\prime})\in\mathcal{P}$. Thus, we have
[…]
Now, we lower bound $\left\lvert\mathcal{P}\right\rvert$. By the inclusion-exclusion principle, we know that $|\mathcal{P}|$ is at least
$|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B}\neq x_{B^{\prime}}\right\}|-|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B}\in B^{\prime}\right\}|-|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B^{\prime}}\in B\right\}|.$
Note that $|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B}\neq x_{B^{\prime}}\right\}|\geqslant|\mathcal{B}_{1}|M$ since there are at least $M$ choices for $B^{\prime}$ once $B$ is chosen (see the definition of $M$). On the other hand, we know that $|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B}\in B^{\prime}\right\}|\leqslant\penalty 10000|\mathcal{B}_{1}|\binom{n-1}{s-1}$ since there are at most $\binom{n-1}{s-1}$ choices for $B^{\prime}$. Similarly, we have $|\left\{(B,B^{\prime})\in\mathcal{B}_{1}\times\mathcal{B}_{1}\mid x_{B^{\prime}}\in B\right\}|\leqslant|\mathcal{B}_{1}|\binom{n-1}{s-1}$. Therefore, we can conclude that
$|\mathcal{P}|\geqslant|\mathcal{B}_{1}|M-2|\mathcal{B}_{1}|{n-1\choose s-1}$.
This completes the proof. ∎
Now, we are ready to upper bound $M$.
Claim 2.10.
We have $M=O_{d}((K+1)n^{s-1})$.
Proof.
If $M\leqslant 2{n-1\choose s-1}$, then the claim is trivially true. Otherwise, combining Claims 2.8 and 2.9, we get
$\left|\mathcal{U}\right|=\Omega_{d}\left(n^{d-s}\left(M-2{n-1\choose s-1}\right)\right)=\Omega_{d}(n^{d-s}M)-O_{d}(n^{d-1}).$
Now by Corollary 2.5, we have
$\left|\mathcal{U}\right|\leqslant\binom{n}{d}-|\mathcal{F}_{1}|\leqslant\binom{n}{d}-|\mathcal{F}|+|\mathcal{F}_{\geqslant 2}|=\binom{n-1}{d}-|\mathcal{F}|+O_{d}(n^{d-1}).$
Combining with the assumption that $|\mathcal{F}|\geqslant{n-1\choose d}-Kn^{d-1}$, we can conclude that
$Kn^{d-1}=\Omega_{d}(n^{d-s}M)-O_{d}(n^{d-1}).$
After rearranging, we get $M=O_{d}((K+1)n^{s-1})$. ∎
Thus by Claim 2.10, we can conclude that
$|\mathcal{F}\triangle\mathcal{F}^{*}|\leqslant Kn^{d-1}+2|\mathcal{F}\setminus\mathcal{F}^{*}|\leqslant Kn^{d-1}+O_{d}(n^{d-1})+O_{d}((K+1)n^{d-1})=O_{d}((K+1)n^{d-1}),$
and Proposition 2.7 follows. ∎
2.3. Stability analysis
We are now ready to perform the stability analysis and finish the proof of Theorem 1.4.
Proof of Theorem 1.4.
Let $\mathcal{F}\subseteq\binom{[n]}{d+1}$ be an $s$-witness family with $1\leqslant s\leqslant d/2$ and suppose $|\mathcal{F}|\geqslant\binom{n-1}{d}$. Then by Proposition 2.7, there exists $x_{0}\in[n]$ such that the star family centered at $x_{0}$ defined as
$\mathcal{F}^{*}=\left\{F^{*}\in\binom{[n]}{d+1}\,\middle|\,x_{0}\in F^{*}\right\}$
satisfying $|\mathcal{F}\triangle\mathcal{F}^{*}|=O_{d}(n^{d-1})$. Let $L=|\mathcal{F}^{*}\setminus\mathcal{F}|$ and note that $L=O_{d}(n^{d-1})$. Our goal is to show that $L=0$.
We will give estimates regarding the following sets:
  1. (1)
    $\mathcal{S}_{1}\overset{\rm{def}}{=}\{B\in\mathcal{B}\mid x_{0}\in B\}$;
  2. (2)
    $\mathcal{S}_{2}\overset{\rm{def}}{=}\{B\in\mathcal{B}_{\geqslant 2}\mid x_{0}\not\in B\}$;
  3. (3)
    $\mathcal{B}_{1}(x_{0})\overset{\rm{def}}{=}\{B\in\mathcal{B}_{1}\mid x_{B}=x_{0}\}$;
  4. (4)
    $\mathcal{B}_{1}^{\prime}\overset{\rm{def}}{=}\{B\in\mathcal{B}_{1}\mid x_{B}\neq x_{0},\,x_{0}\notin B\}$.
Claim 2.11.
We have
$\sum_{B\in\mathcal{S}_{1}}|\mathcal{F}_{B}|=O_{d}(L/n).$
Proof.
Fix $B\in\mathcal{S}_{1}$. We first lower bound the number of sets $F\in\mathcal{F}^{*}\setminus\mathcal{F}$ with $B\subseteq\mathcal{F}$. For any $A\in\mathcal{A}_{B}$, we know that the number of sets $F\in\mathcal{F}^{*}$ that contain $A\cup B$ is at most $\binom{n-s-\left\lvert A\right\rvert}{d+1-s-\left\lvert A\right\rvert}=O_{d}(n^{d-s})$. On the other hand, there are $\binom{n-s}{d+1-s}$ sets in $\mathcal{F}^{*}$ that contain $B$. Since any set $F\in\mathcal{F}^{*}\cap\mathcal{F}$ with $B\subseteq\mathcal{F}$ must contain $A\cup B$ for some $A\in\mathcal{A}_{B}$, we know that the number of $F\in\mathcal{F}^{*}\setminus\mathcal{F}$ with $B\subseteq\mathcal{F}$ is at least $\binom{n-s}{d-s+1}-O_{d}(n^{d-s})\left\lvert\mathcal{A}_{B}\right\rvert=\Omega_{d}(n^{d-s+1})$.
Since every $F^{*}\in\mathcal{F}\setminus\mathcal{F}$ contains at most $d\choose s-1$} sets $B\in\mathcal{S}_{1}$, we obtain
$L=|\mathcal{F}^{*}\setminus\mathcal{F}|\geqslant\binom{d}{s-1}^{-1}\cdot\Omega_{d}(n^{d-s+1})\cdot|\mathcal{S}_{1}|.$
Thus, we have $|\mathcal{S}_{1}|=O_{d}(Ln^{s-d-1})$. On the other hand, by Claim 2.3, for any $B\in\mathcal{S}_{1}$, we know that $|\mathcal{F}_{B}|=O_{d}(n^{d+1-s-\alpha_{B}})=O_{d}(n^{d-s})$, so
$\sum_{B\in\mathcal{S}_{1}}|\mathcal{F}_{B}|=O_{d}(n^{d-s})\cdot|\mathcal{S}_{1}|=O_{d}(L/n).\qed$
Claim 2.12.
We have
$\sum_{B\in\mathcal{S}_{2}}|\mathcal{F}_{B}|=O_{d}(L/n).$
Proof.
The proof is similar to the one of Claim 2.11. Fix $B\in\mathcal{S}_{2}$. We first lower bound the number of sets $F\in\mathcal{F}^{*}\setminus\mathcal{F}$ with $B\subseteq\mathcal{F}$. For any $A\in\mathcal{A}_{B}$, we know that $\left\lvert A\right\rvert\geqslant 2$ since $B\in\mathcal{B}_{\geqslant 2}$. By a direct calculation, the number of sets $F\in\mathcal{F}^{*}$ that contains $A\cup B$ is at most $\binom{n-s-1-\left\lvert A\right\rvert}{d+1-s-\left\lvert A\right\rvert}=O_{d}(n^{d-s-1})$ (equality is achieved when $x_{0}\in A$ and $\left\lvert A\right\rvert=2$). On the other hand, there are $\binom{n-s-1}{d-s}$ sets in $\mathcal{F}^{*}$ that contains $B$. Since any set $F\in\mathcal{F}^{*}\cap\mathcal{F}$ with $B\subseteq\mathcal{F}$ must contain $A\cup B$ for some $A\in\mathcal{A}_{B}$, we know that the number of $F\in\mathcal{F}^{*}\setminus\mathcal{F}$ with $B\subseteq\mathcal{F}$ is at least $\binom{n-s-1}{d-s}-O_{d}(n^{d-s-1})\left\lvert\mathcal{A}_{B}\right\rvert=\Omega_{d}(n^{d-s})$.
Since every $F^{*}\in\mathcal{F}\setminus\mathcal{F}$ contains at most $d+1\choose s$} sets $B\in\mathcal{S}_{2}$, we obtain
$L=|\mathcal{F}^{*}\setminus\mathcal{F}|\geqslant\binom{d+1}{s}^{-1}\cdot\Omega_{d}(n^{d-s})\cdot|\mathcal{S}_{2}|.$
Thus, we have $|\mathcal{S}_{2}|=O_{d}(Ln^{s-d})$. On the other hand, by Claim 2.3, for any $B\in\mathcal{S}_{2}$, we know that $|\mathcal{F}_{B}|=O_{d}(n^{d+1-s-\alpha_{B}})=O_{d}(n^{d-s-1})$, so
$\sum_{B\in\mathcal{S}_{2}}|\mathcal{F}_{B}|=O_{d}(n^{d-s-1})\cdot|\mathcal{S}_{2}|=O_{d}(L/n).\qed$
Since we are focusing on the sets $F\in\mathcal{F}$ that are not captured by $\mathcal{F}^{*}$, we need restrict our injection to the sets in $n]\setminus\{x_{0}\}$. Thus, we define
$\mathcal{F}_{1}(x_{0})\overset{\rm{def}}{=}\bigcup_{B\in\mathcal{B}_{1}(x_{0})}\mathcal{F}_{B},\quad\quad$ $\mathcal{F}_{1}^{\prime}\overset{\rm{def}}{=}$ $\bigcup_{B\in\mathcal{B}_{1}^{\prime}}\mathcal{F}_{B},$
$\mathcal{P}^{\prime}\overset{\rm{def}}{=}\mathcal{P}\cap(\mathcal{B}_{1}(x_{0})\times\mathcal{B}_{1}^{\prime}),$ $\quad\mathcal{U}^{\prime}\overset{\rm{def}}{=}$ $\bigcup_{(B,B^{\prime})\in\mathcal{P}^{\prime}}\mathcal{U}(B,B^{\prime}).$
Note that $\mathcal{U}^{\prime}\subseteq\binom{[n]\setminus\{x_{0}\}}{d}$ since any set $E\in\mathcal{U}(B,B^{\prime})$ does not contain $x_{0}=x_{B}$ if $B\in\mathcal{B}_{1}(x_{0})$. Thus, the injection $\varphi$ we defined right after Lemma 2.4 gives the following corollary.
Corollary 2.13.
We have
$\left\lvert\mathcal{F}_{1}(x_{0})\right\rvert+\left\lvert\mathcal{F}_{1}^{\prime}\setminus\mathcal{F}^{*}\right\rvert+\left\lvert\mathcal{U}^{\prime}\right\rvert\leqslant\binom{n-1}{d}$.
Proof.
Note that $\varphi(\mathcal{F}_{1}(x_{0}))$, $\varphi(\mathcal{F}_{1}^{\prime}\setminus\mathcal{F}^{*})$, and $\mathcal{U}^{\prime}$ are set families in $\binom{[n]\setminus\{x_{0}\}}{d}$ and they are pairwise disjoint by Lemma 2.4. ∎
We also need the following variant of Claim 2.9.
Lemma 2.14.
[…]
Proof.
[…]
Note that for each fixed $B^{\prime}\in\mathcal{B}_{1}^{\prime}$, there are at most $\binom{n-2}{s-1}$ possible $B\in\mathcal{B}_{1}(x_{0})$ such that $x_{B^{\prime}}\in B$. Thus, there are at least $\left\lvert\mathcal{B}_{1}(x_{0})\right\rvert-\binom{n-2}{s-1}$ pairs of the form $(B,B^{\prime})\in\mathcal{P}^{\prime}$ for each $B^{\prime}\in\mathcal{B}_{1}^{\prime}$. Recall that we defined $M\geqslant 0$ such that $|\mathcal{B}_{1}(x_{0})|=|\mathcal{B}_{1}|-M$. From Claims 2.8 and 2.10 with $K=0$, we have
$\left\lvert\mathcal{B}_{1}(x_{0})\right\rvert-\binom{n-2}{s-1}=\left\lvert\mathcal{B}_{1}\right\rvert-M-\binom{n-2}{s-1}=\Omega_{d}(n^{s}).$
[…]
Now we are ready to finish the proof of Theorem 1.4. Note that $\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\mathcal{B}_{1}(x_{0})\cup\mathcal{B}_{1}^{\prime}=\mathcal{B}$, and hence
$\left\lvert\mathcal{F}\right\rvert\leqslant\sum_{B\in\mathcal{S}_{1}}|\mathcal{F}_{B}|+\sum_{B\in\mathcal{S}_{2}}|\mathcal{F}_{B}|+\left\lvert\mathcal{F}_{1}(x_{0})\right\rvert+\left\lvert\mathcal{F}_{1}^{\prime}\right\rvert$.
[…]
$\left\lvert\mathcal{F}\right\rvert$ $\leqslant O_{d}(L/n)+\left\lvert\mathcal{F}_{1}(x_{0})\right\rvert+\left\lvert\mathcal{F}_{1}^{\prime}\setminus\mathcal{F}^{*}\right\rvert+\left\lvert\mathcal{F}_{1}^{\prime}\cap\mathcal{F}^{*}\right\rvert$
$\leqslant O_{d}(L/n)+\binom{n-1}{d}-\left\lvert\mathcal{U}^{\prime}\right\rvert+\left\lvert\mathcal{F}_{1}^{\prime}\cap\mathcal{F}^{*}\right\rvert$. (2)
For any fixed $B\in\mathcal{B}_{1}^{\prime}$, we know that $|\mathcal{F}_{B}\cap\mathcal{F}^{*}|\leqslant\binom{n-s-2}{d-s-1}$ since any set $F\in\mathcal{F}_{B}\cap\mathcal{F}^{*}$ must contain $B\cup\{x_{0},x_{B}\}$. Therefore, we have
[…]
Combining Lemmas 2.14, 2 and 3, we have
[…]
Recall that we assumed $\left\lvert\mathcal{F}\right\rvert\geqslant\binom{n-1}{d}$. It follows that $\left\lvert\mathcal{B}_{1}^{\prime}\right\rvert=O_{d}(n^{s-d-1}L)$
Now, we bound $\left\lvert\mathcal{F}\setminus\mathcal{F}^{*}\right\rvert$. Note that if $F\in\mathcal{F}\setminus\mathcal{F}^{*}$, then $F\in\mathcal{F}_{B}$ for some $B\in\mathcal{S}_{2}\cup\mathcal{B}_{1}^{\prime}$. Therefore,
$\left\lvert\mathcal{F}\setminus\mathcal{F}^{*}\right\rvert\leqslant\sum_{B\in\mathcal{S}_{2}}|\mathcal{F}_{B}|+\left\lvert\mathcal{F}_{1}^{\prime}\setminus\mathcal{F}^{*}\right\rvert\leqslant O_{d}(L/n)+O_{d}(n^{d-s}\left\lvert\mathcal{B}_{1}^{\prime}\right\rvert)=O_{d}(L/n),$
where the second inequality follows from Claims 2.12 and 2.3.
However, we have
$\left\lvert\mathcal{F}\setminus\mathcal{F}^{*}\right\rvert=\left\lvert\mathcal{F}\right\rvert-\left\lvert\mathcal{F}^{*}\right\rvert+\left\lvert\mathcal{F}^{*}\setminus\mathcal{F}\right\rvert\geqslant\binom{n-1}{d}-\binom{n-1}{d}+L=L.$
Therefore, $O_{d}(L/n)\geqslant L$, and hence $L=\left\lvert\mathcal{F}^{*}\setminus\mathcal{F}\right\rvert=0$ when $n$ is sufficiently large. This implies that $\mathcal{F}^{*}\subseteq\mathcal{F}$, and it is clear that $\mathcal{F}^{*}$ is a maximal $s$-witness family. This completes the proof. ∎
3. Non-star constructions
We first prove Theorem 1.5 in Section 3.1, and then prove Theorem 1.6 in Section 3.2
3.1. A non-star construction for $d/2<s\leqslant d-1$
Now we give a construction that proves Theorem 1.5.
Proof of Theorem 1.5.
Fix any integer $m\in[s+2,n-1]$. Consider the set family $\mathcal{F}\overset{\rm{def}}{=}\mathcal{A}_{1}\cup\mathcal{A}_{2}\cup\mathcal{A}_{12}$ defined as follows:
$\mathcal{A}_{1}\overset{\rm{def}}{=}$ $\left\{\{1\}\cup A\,\middle|\,A\in\binom{[3,n]}{d},\,\left\lvert A\cap[3,m]\right\rvert\geqslant s\right\},$
$\mathcal{A}_{2}\overset{\rm{def}}{=}$ $\left\{\{2\}\cup A\,\middle|\,A\in\binom{[3,n]}{d},\,\left\lvert A\cap[3,m]\right\rvert<s\right\},$
$\mathcal{A}_{12}\overset{\rm{def}}{=}$ $\left\{\{1,2\}\cup A\,\middle|\,A\in\binom{[3,n]}{d-1}\right\}$,
It is clear that $\mathcal{F}$ is not a star. Moreover, we have
$|\mathcal{F}|=\binom{n-2}{d}+\binom{n-2}{d-1}=\binom{n-1}{d}$.
We choose the witness $B_{F}$ for each $F\in\mathcal{F}$ as follows.
  • $F\in\mathcal{A}_{12}$: $F=\{1,2\}\cup A$ for some $A\in\binom{[3,n]}{d-1}$. We set $B_{F}\subseteq A$ to be any arbitrary subset of size $s$.
  • $F\in\mathcal{A}_{1}$: $F=\{1\}\cup A$ for some $A\in\binom{[3,n]}{d}$ such that $|A\cap[3,m]|\geqslant s$. We set $B_{F}\subseteq A\cap[3,m]$ to be any arbitrary subset of size $s$.
  • $F\in\mathcal{A}_{2}$: $F=\{2\}\cup A$ for some $A\in\binom{[3,n]}{d}$ such that $|A\cap[3,m]|<s$. We set $B_{F}\subseteq A$ to be any subset of size $s$ such that $|B_{F}\cap[m+1,n]|\geqslant d-s+1$. This is well-defined by the assumption that $|A\cap[m+1,n]|\geqslant d-|A\cap[3,m]|\geqslant d-s+1$ and $s\geqslant d-s+1$.
Now we justify that the above constructions are indeed valid witnesses.
  • $F\in\mathcal{A}_{12}$: Since $\{1,2\}\subseteq F$ and $\{1,2\}\cap F^{\prime}\neq\varnothing$ for all $F^{\prime}\in\mathcal{F}$, so we have $\{1,2\}\cap(F\cap F^{\prime})\neq\varnothing$. Since $B_{F}\cap\{1,2\}=\varnothing$, we have $B_{F}$ is a valid witness for $F$.
  • $F\in\mathcal{A}_{1}$: Since $1\in F$ and $1\not\in B_{F}$, any $F^{\prime}\in\mathcal{F}$ satisfying $F\cap F^{\prime}=B_{F}$ must be in $\mathcal{A}_{2}$. However, since $|F^{\prime}\cap[3,m]|<s$ and $B_{F}\subseteq A\cap[3,m]$, we cannot have $F^{\prime}\cap F=B_{F}$.
  • $F\in\mathcal{A}_{2}$: Similarly, since $2\in F$ and $2\not\in B_{F}$, any $F^{\prime}\in\mathcal{F}$ satisfying $F\cap F^{\prime}=B_{F}$ must be in $\mathcal{A}_{1}$. Then we must have $\left\lvert F^{\prime}\cap[m+1,n]\right\rvert\leqslant d-s$. Since we have $B_{F}\supseteq A\cap[m+1,n]$ and $|A\cap[m+1,n]\geqslant d-s+1$, there must exist an element in $B_{F}$ that is not in $F^{\prime}$, so $F\cap F^{\prime}\neq B_{F}$.
This shows that $\mathcal{F}$ is an $s$-witness family. ∎
3.2. A non-star construction with arbitrarily large piercing number
The construction for Theorem 1.6 is similar to the one given in Section 3.1. In particular, when the piercing number $k=2$, the following construction recovers the one in Section 3.1 with $m=s+2$.
Proof of Theorem 1.6.
Let $U=[n]\setminus[k]=\{k+1,\ldots,n\}$. Let $S_{1},S_{2},\ldots,S_{k-1}\subseteq U$ be a collection of pairwise disjoint subsets of size $s$. This can be done provided that $n\geqslant k+(k-1)s$. For $i\in[k-1]$ and $m\geqslant s$, let us define the families
$\mathcal{A}_{i}(m)\overset{\rm{def}}{=}\left\{A\in{U\choose m}\,\middle|\,S_{i}\subseteq A\right\},$
and
$\mathcal{A}_{k}(m)\overset{\rm{def}}{=}\left\{A\in{U\choose m}\,\middle|\,S_{i}\not\subseteq A~\forall i=1,\ldots,k-1\right\}.$
Using the assumption that $2s>d$, we know that these families are pairwise disjoint and they cover the sets in $U\choose m$} if $m\leqslant d$. For a non-empty subset $I\subseteq[k]$, we define
$\mathcal{A}_{I}\overset{\rm{def}}{=}\bigsqcup_{i\in I}\mathcal{A}_{i}\left(d+1-|I|\right)\subseteq{U\choose d+1-|I|}$.
Let $\mathcal{F}_{I}\overset{\rm{def}}{=}\{I\cup A\mid A\in\mathcal{A}_{I}\}$ and define $\mathcal{F}\overset{\rm{def}}{=}\bigsqcup_{\varnothing\neq I\subseteq[k]}\mathcal{F}_{I}$. We claim that $\mathcal{F}$ is an $s$-witness family of size $n-1\choose d$} with piercing number $k$. First, let us compute the size of $\mathcal{F}$. Note that for each $t=1,\ldots,k$, we have
$\sum_{I\in{[k]\choose t}}|\mathcal{A}_{I}|=\sum_{i=1}^{k}{k-1\choose t-1}|\mathcal{A}_{i}|={k-1\choose t-1}{n-k\choose d+1-t}$,
and so
$|\mathcal{F}|=\sum_{\varnothing\neq I\subseteq[k]}|\mathcal{A}_{I}|=\sum_{t=1}^{k}{k-1\choose t-1}{n-k\choose d+1-t}={n-1\choose d}$.
Next, we show that $\mathcal{F}$ is indeed an $s$-witness family. We begin by defining the witness sets $B_{F}$ for each $F\in\mathcal{F}$. Suppose that $F\in\mathcal{F}_{I}$ with nonempty $I\subseteq[k]$, so that there exists $A\in\mathcal{A}_{I}$ such that $F=I\cup A$. Then we define $B_{F}$ in the following way:
  • If $S_{i}\subseteq A$ for some $i\in[k-1]$, then we set $B_{F}=S_{i}$.
  • Otherwise, let $J$ be the set of indices $j\in[k-1]\setminus I$ such that $A\cap S_{j}\neq\varnothing$. For every $j\in J$, fix an arbitrary element $x_{j}\in S_{j}\cap A$. Let $B_{F}$ be an arbitrary subset of $A\setminus\{x_{j}\mid j\in J\}$ of size $s$.
Note that in the second case we have $|J|\leqslant k-|I|$. Thus, we have
$|A\setminus\{x_{j}\mid j\in J\}|=|A|-|J|=d+1-|I|-|J|\geqslant d+1-k\geqslant s,$
and hence it is possible to pick an $s$-subset $B_{F}\subseteq A\setminus\{x_{j}\mid j\in J\}$. Furthermore, we note that $B_{F}\subseteq U$ for all $F\in\mathcal{F}$.
Now, we check the $s$-witness condition. That is, we check that $F\cap F^{\prime}\neq B_{F}$ for all $F,F^{\prime}\in\mathcal{F}$. Suppose that $F=I\cup A$ and $F^{\prime}=I^{\prime}\cup A^{\prime}$ with nonempty $I,I^{\prime}\subseteq[k]$, $A\in\mathcal{A}_{I}$, and $A^{\prime}\in\mathcal{A}_{I^{\prime}}$. We check the condition in the following case analysis.
  • If $I\cap I^{\prime}\neq\varnothing$, then $F\cap F^{\prime}\supseteq I\cap I^{\prime}$. Since $I\cap I^{\prime}$ is disjoint from from $B_{F}\subseteq U$, we have $F\cap F^{\prime}\neq B_{F}$ in this case.
  • If $I\cap I^{\prime}=\varnothing$ and $S_{i}\subseteq A$ for some $i\in[k-1]$, then $i\in I$ and $B_{F}=S_{i}$. It follows that $i\not\in I^{\prime}$ and hence $B_{F}=S_{i}\not\subseteq A^{\prime}$.
  • Otherwise, we have $I\cap I^{\prime}=\varnothing$ and $S_{i}\not\subseteq A$ for all $i\in[k-1]$. It follows that $k\in I$. Since $A^{\prime}\in\mathcal{F}_{I^{\prime}}$ and $I^{\prime}\subseteq[k-1]$, there exists $i^{\prime}\in I^{\prime}$ such that $S_{i^{\prime}}\subseteq A^{\prime}$.
    • If $A\cap S_{i^{\prime}}=\varnothing$, then we have
      $|F\cap F^{\prime}|=|A\cap A^{\prime}|=|A\cap(A^{\prime}\setminus S_{i^{\prime}})|\leqslant|A^{\prime}|-|S_{i^{\prime}}|\leqslant d-s<s,$
      and hence $F\cap F^{\prime}\neq B_{F}$.
    • Otherwise, we have $A\cap S_{i^{\prime}}\neq\varnothing$. By construction there is an element $x_{i^{\prime}}\in A\cap S_{i^{\prime}}$ which does not lie in $B_{F}$. On the other hand, since $S_{i^{\prime}}\subseteq A^{\prime}$, we clearly have $x_{i^{\prime}}\in A\cap A^{\prime}\subseteq F\cap F^{\prime}$. Therefore, $F\cap F^{\prime}\neq B_{F}$.
By the case analysis above, we know that $\mathcal{F}$ is an $s$-witness family.
It remains to show that $\tau(\mathcal{F})=k$. It is clear that $[k]$ pierces every set in $\mathcal{F}$, and hence $\tau(\mathcal{F})\leqslant k$. Now, assume for the sake of contradiction that there is some (k1)(k-1)-set $T=I\sqcup R$, where $I\subseteq[k]$ and $R\subseteq U$, that intersects every set in $\mathcal{F}$. For any element $j\in[k]\setminus I$, we know that $R$ intersects every set in $\mathcal{A}_{j}(d)$. For each $j\in[k-1]\setminus I$, then we must have $R\cap S_{j}\neq\varnothing$. Otherwise, we can find a set $A\subseteq U\setminus R$ of size $d$ such that $S_{j}\subseteq A$ and $A\cap R=\varnothing$. This can be done since $\left\lvert U\setminus R\right\rvert=\left\lvert U\right\rvert-\left\lvert R\right\rvert\geqslant n-k-(k-1)\geqslant d$. Therefore, we know that $\left\lvert(I\cup R)\cap S_{j}\right\rvert=1$ for all $j\in[k-1]$, and $k\notin I$. Thus, $R$ must intersect every set in $\mathcal{A}_{k}(d)$. However, we can find a set $A\subseteq U\setminus R$ of size $d$ that does not contain any $S_{i}$ with $i\in[k-1]$ as a subset. This is because we only need to avoid the set $R$ and also avoid one element from $S_{i}$ for each $i\in I$. This is doable since $\left\lvert U\right\rvert-\left\lvert R\right\rvert-\left\lvert I\right\rvert=n-k-(k-1)\geqslant d$.
We can conclude that $\mathcal{F}$ is an $s$-witness family of size $\binom{n-1}{d}$ and it has piercing number $k$. This completes the proof. ∎